OpenStudy (anonymous):

int e^-x/(9-e^-x) dx shouldn't be the -ln(9-e^-x) ? whereas both text & maple shows ln(9-e^-x), why?

OpenStudy (anonymous):

no, because if you did u-sub with u=9-e^-x, du=e^-x dx

OpenStudy (anonymous):

i got the second answer you posted.

OpenStudy (anonymous):

int e^-x/(9-e^-x) dx =int 1/(9-e^-x) d(9-e^-x) <---- note that d(-e^-x) = -(-e^-x) dx=e^-x dx =ln (9-e^-x) +

OpenStudy (anonymous):

u = 9 - e^-x => du = e^-x dx => Int ( du/ u ) = ln ( 9 - e^-x) + C

OpenStudy (anonymous):

ohh gosh I was doing the wrong substitution instead of u=denominator i was doing u=numerator

OpenStudy (anonymous):

thanks @ all